Gun shoots a bullet at 2500m/s at an angle of 65 degrees above the horizontal. Neglecting air resistance, find its range.
1
Expert's answer
2019-10-22T10:03:37-0400

The range


"R=\frac{v_0^2\sin 2\theta}{g}=\frac{2500^2\sin (2\times 65^{\circ})}{9.8}=488549\: \rm m"